# Pipeline Parallelism with Controllable Memory

### v1

@@ 2 How to build a Pipeline
Repeating: More minibatches are then introduced. The building blocks are repeated and woven together to form a pipeline. In Figure 1 (top), the repeating building blocks are shown in different shades of grey. Notably, legit building blocks are required to repeat without a collision, namely, the passes from two building blocks should not overlap with each other.
@@
Figure 1: How to build a Pipeline
Most of the existing pipeline schedules can be explained under this framework. Besides the 1F1B and Eager 1F1B (Zhuang et al., 2023) shown in Figure 1, we show the interleaved 1F1B (Shoeybi et al., 2019), ZB-H1 (Qi et al., 2023) and a series of well-known pipeline in a more extensive gallery (see Appendix G).
@@ 2.2 Calculating the Peak Memory
Two quantities are crucial for the calculation of peak memory, the lifespan of a stage, and the repeating interval of the building blocks, both of which are illustrated in Figure 1. The lifespan of a stage is defined as the amount of time in between the starting of the F pass and the ending of B or W pass. A piece of activation memory is allocated at the starting of F, and retained in the RAM throughout the lifespan until it is consumed by both B and W. The peak memory consumption can be calculated by finding the maximum number of minibatches whose lifespans overlap with that of every other minibatch. Using $l$ to denote lifespan, $T$ to denote the repeating interval and $m$ the size of activation memory for a single microbatch, we have the relation.
@@ 3 Memory Efficient Building Blocks
With the above framework, we can conveniently analyse the memory consumption pattern of existing pipelines. To our knowledge, all existing pipelines are memory inefficient due to the imbalanced memory usage. The imbalanced memory comes from the innate heterogeneity of the lifespans across stages. From Figure 1(a), we can easily see that the lifespan of the stages differs greatly from each other, with the first stage having the longest lifespan. Consequently, it causes a memory bottleneck on the first device and under utilization of memory on all other devices. To resolve this problem, we introduce a family of novel building blocks which we refer to as V-Shape building blocks. The insight of the V-Shape building block comes from Equation 1 which says that the peak memory depends on the sum of the lifespans. Therefore, when we place multiple stages on the same device, we should always collocate stages of long lifespans with those of short lifespans. When the total sum of lifespans is fixed, balanced placement always means higher memory efficiency. This can be demonstrated by Figure 2, the parallel schedule (used in interleaved 1F1B) is imbalanced and has a memory bottleneck proportional to $l_{1}+l_{4}$, while in the V-Shape schedule it is $l_{1}+l_{6}$.
The V-Shape schedule requests the model to be partitioned into stages twice the number of devices and the device placement of the second half of stages to be in reverse order as the first half. We can then further control the offsets between the passes to generate diverse building blocks. As the offsets directly determine the lifespan of each stage and therefore the peak memory by Equation 1, in the following, we introduce different V-Shape building blocks by their peak memory usage.
@@
3.1 V-Shape Building Blocks
Intuitively, in pipeline parallelism there is usually a trade-off between bubble rate and memory consumption (Qi et al., 2023), because we need more memory to fill in the warm-up and cool-down bubbles. Following this intuition, we come up with 3 novel V-Shape building blocks with different memory consumption and bubble rate, controlled by different offsets in Figure 3. The building block of V-Min shown in Figure 3(c) has the minimum offsets thus the minimum memory consumption. Pushing to extreme throughput, V-ZB shown in Figure 3(d) can reduce bubble rate to 0. The building block of V-Half sits between the two extremes and consumes half of the activation memory needed by 1F1B. Note that although we show only 3 cases here, we can have a flexible memory consumption using the search algorithm illustrated in Appendix A.
We assume the model is uniformly partitioned, and use $m$ to represent the activation memory of a single microbatch for each stage. We denote $d$ as the number of devices, $M$ as the total activation memory of the entire model for a single microbatch, $l_{max}=\max_{i}{(\sum_{s\in S_{i}}{l^{s}})}$ as the maximum total lifespan across all devices, we can analyse the lifespan and peak activation memory of 1F1B and V-Shape building blocks, shown in Table 1. Note that we ignored the constant terms when measuring $l_{max}$ for simplicity in the analysis, e.g. accurately for V-ZB $l_{max}=12d-2$ but we use $l_{max}=12d$ instead.
@@ 3.1 V-Shape Building Blocks
Table 1: Activation Memory (Asymptotic) of V-Shape Building Blocks Compared to 1F1B
Building Block Interval $T$ $l_{max}$ $m$ Peak Memory
[…]
Notably, to avoid collision the offsets in building block for V-Min and V-Half is slightly different on different $d$. The details are in Appendix F.
3.2 V-Shape Pipeline Schedules
Based on the framework in Section 2 and our V-Shape building blocks, we present the final schedules of V-Min, V-Half and V-ZB in Figure 4. Assuming F B W have equal run time, V-Min reduces the bubble rate to 2/3 of 1F1b, V-Half to 1/2 and while V-ZB achieves zero bubble similar to (Qi et al., 2023).
@@ 3.2 V-Shape Pipeline Schedules
In real cases where F, B and W may have different run times, V-Min suffers from a repeating bubble. As shown in Figure 5, there exists a bubble for every repeating interval $T$. As a result, the bubble grows with respect to the number of microbatches. Although V-Half may also encounter the same issue, we surprisingly find that it generates patterns that tessellates well in most empirical cases. As illustrated in Figure 5(b), the throughput of V-Half is robust to the variation of run times. Additionally, we show that the bubbles of V-ZB will never grow when increasing the number of microbatches. We leave the related discussions in Appendix E.
@@
Figure 5: Repeating Bubble of V-Min compared to V-Half. We take a repeating $d\times T$ grid from both schedule and assign F B W with different values. The result shows V-Min has a bubble for every repeating grid, while V-Half does not.
@@ 3.3 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks (see Appendix I), to prove the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule with the same bubble rate as Interleaved 1F1B(Narayanan et al., 2021) but lower memory consumption.
@@ 4 Experiments
We construct our experiments to show three conclusions: a) The throughput and memory of V-Min, V-Half and V-Min aligns with the theoretical analysis in Section 3.2; b) Memory-saving methods including V-Min and V-ZB can bring acceleration; c) Our methods still perform best when combined with other state-of-the-art techniques.
@@ 4.1 Setup
We evaluate our methods using a serious of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 4 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stages to compensate for the embedding or lm-head layers, which can otherwise become the bottleneck for the pipeline and interfere the efficiency.
Table 2: Models used in experiments
@@
Our experiments majorly focuses on the following pipeline parallel schedules: 1) V-Min, V-Half and V-ZB: Schedules introduced in Section 3.2; 2) 1F1B and Interleaved 1F1B: Method implemented in Megatron-LM; 3) 1F1B-R: 1F1B used with full activation rematerialization; 4) ZB-1P and ZB-2P: The adaptive zero-bubble method introduced in (Qi et al., 2023) with activation memory limit set to the 1x/2x times of 1F1B.
@@ 4.2 Comparing Pipeline Schedules
In Figure 6 we present a comparison of the throughput measured in FLOPS utilization (MFU) and activation memory consumption of each pipeline schedule using different settings, we find V-ZB outperforms all other methods, which aligns with Figure 4. When comparing the activation memory consumption, V-Min and V-Half stand out by significantly reducing activation memory to approximately 1/3 and 1/2, while other methods’ memory is similar. More details of our experiments and definition of metrics can be found in Appendix D.1.
Notably we find V-Min has a comparable throughput against 1F1B, but its throughput falls behind 1F1B at a larger number of microbatch due to the aforementioned bubble in Figure 5(a), as discussed in 3.2. However, it still outperforms 1F1B with full activation rematerialization, providing a strong alternative for saving memory.
@@ 4.3 When to Save Memory
Another benefit of saving memory is that we can potentially use the extra memory for an increased microbatch size, which leads to a higher arithmetic intensity. We present the result in Figure 8. On bigger models, where memory pressure is higher and hence microbatch size is smaller, V-Half schedule can surpass V-ZB and other baselines because of its arithmetic intensity gain. This observation does not apply for V-Min, implying its arithmetic intensity gain can not compensate for the increased bubble. Doubling/Tripling the microbatch size for V-Half/V-Min results in a slightly higher activation memory than the other methods. This reflects the constant factors we ignored in Section 3.2. The increase is less significant as the number of devices grows.
@@ 4.4 Combining with Existing Techniques
We present our method in the context of LLM training where it is used with various other techniques. The following techniques are considered: a) Flash Attention (Dao et al., 2022; Dao, 2023); b) Tensor Parallel (Narayanan et al., 2021) and Sequence Parallel (Korthikanti et al., 2023); c) Distributed Optimizer provided in Megatron-LM; The implementations are all from Megatron-LM (Narayanan et al., 2021). Both our methods and the baseline methods are combined with the above techniques. Similar to the evaluation method in (Kim et al., 2023), we perform a grid search on the following parameters: the size of PP; the size of TP; the size of DP; the microbatch size ($mb$). We use 40 GPUs in this experiment. For each method, the best result from the grid search is reported.
We present the best result for each pipeline parallel schedule in table 6 and the corresponding parameters. We find that when sequence length is smaller and hence the memory budget is more abundant, V-ZB performs the best due to the elimination of bubble. When we increase the sequence length and the number of microbatches, V-Half performs the best because of its memory efficiency. The detailed data and analysis of grid search can be found in the Appendix D.3.
@@ 5 Conclusion And Future Work
In this work, we present a framework that designs pipeline parallel schedules by focusing on their repeating building blocks. We introduce a method to calculate peak activation memory from the building block. Based on this, we design a family of memory-efficient building blocks. We discuss and evaluate three representative methods from the family, namely V-Min, V-Half and V-ZB. We demonstrate with experiments that methods from this family advance the pareto frontier of throughput and memory in large model training, thereby accelerating the training of large models in practice. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules in future works.
In the future, we plan to further explore more memory efficient pipeline schedules based on our framework. A prominent limitation of V-Min is that, it suffers from growing bubbles when increasing the number of microbatches. Although V-Half mitigates this issue, there is still a space to further reduce the memory consumption. Using continuous offsets or finer-granularity discretization is a promising way to solve it. We leave it in our future work.
@@ Appendix I Other Memory Efficient Building Blocks
(b) Zero Buble Schedule with 2/3 1F1B memory

### v2

@@ 2 How to Build a Pipeline
Repeating: More microbatches are then introduced. The building blocks are repeated and woven together to form a pipeline. In Figure 1 (top), the repeating building blocks are shown in different shades of grey. Notably, legit building blocks are required to repeat without a collision, namely, the passes from two building blocks should not overlap with each other.
@@
Figure 1: How to build a Pipeline.
Most of existing pipeline schedules can be explained under this framework. Besides the 1F1B and eager 1F1B shown in Figure 1, we show the interleaved 1F1B (Shoeybi et al., 2019), ZB-H1 (Qi et al., 2023) and a series of well-known pipelines in a more extensive gallery (see Appendix G).
@@ 2.2 Calculating the Peak Memory
Two quantities are crucial for the calculation of peak memory, the lifespan of a stage, and the repeating interval of the building blocks, both of which are illustrated in Figure 1. The lifespan of a stage is defined as the amount of time in between the starting of the F pass and the ending of B or W pass. A piece of activation memory is allocated at the starting of F, and retained in the RAM throughout the lifespan until it is consumed by both B and W. The peak memory consumption can be calculated by finding the maximum number of microbatches whose lifespans overlap with that of every other microbatch. Using $l$ to denote lifespan, $T$ to denote the repeating interval and $m$ the size of activation memory for a single microbatch, we have the relation.
@@ 3 Memory Efficient Building Blocks
With the above framework, we can conveniently analyse the memory consumption pattern of existing pipelines. To our knowledge, all existing pipelines are memory inefficient due to two primary reasons: redundant dependency chain, and imbalanced memory usage. Before Zero Bubble (Qi et al., 2023), the backward is often regarded as a single pass, resulting in unnecessarily longer lifespan thus more memory footprint. In this paper, we leverage the backward splitting strategy to remove these redundant lifespan. The imbalanced memory comes from the innate heterogeneity of the lifespans across stages. From Figure 1(a), we can easily see that the lifespan of the stages differs greatly from each other, with the first stage having the longest lifespan. Consequently, it causes a memory bottleneck on the first device and under utilization of memory on all other devices. To resolve this problem, we introduce a family of novel building blocks which we refer to as V-Shape building blocks. The core insight comes from Equation 1 which says that the peak memory depends on the sum of the lifespans. Therefore, when we place multiple stages on the same device, we should always collocate stages of long lifespans with those of short lifespans. When the total sum of lifespans is fixed, balanced placement always means higher memory efficiency. This can be demonstrated by Figure 2, the parallel schedule (used in interleaved 1F1B) is imbalanced and has a memory bottleneck proportional to $l_{1}+l_{4}$, while in the V-Shape schedule it is $l_{1}+l_{6}$.
The V-Shape schedule requests the model to be partitioned into stages twice the number of devices and the device placement of the second half of stages to be in reverse order as the first half. As the offsets directly determine the lifespan of each stage and therefore the peak memory by Equation 1, we can then further control the offsets between passes to generate building blocks with diverse memory.
@@
3.1 Controllable Balanced Memory
In this section, we introduce a simple approach to control the peak memory by using uniform offsets within each half of F and B passes. Note that we are aiming at the memory efficiency, so we are only interested in the building blocks with balanced peak memory across devices.
@@ 3.1 Controllable Balanced Memory
We assume the model is uniformly partitioned, namely, both the computation and memory of each stage is the same. For a single microbatch, we denote the activation memory of each stage as $m$, and the total activation memory of the entire model as $M$. Note that $M=2dm$, where $d$ is the number of devices. To reduce the search space, we use uniform offsets within each half of F and B passes. Specifically, we apply the same offset (denoted as $\delta_{F}^{0}$) between two adjacent F passes within the first $d$ stages. For examples, we set $\delta_{F}^{0}$ as $2$ in Figure 3(b), as $1$ in Figure 3(c) and as $4$ in Figure 3(d). We add similar constraints for the other half of F passes and both half of B passes, and use $\delta_{F}^{1},\delta_{B}^{0},\delta_{B}^{1}$ to denote the offset between adjacent F passes within the last $d$ stages, the offset between adjacent B passes within the first $d$ stages, the offset between adjacent B passes within the last $d$ stages, respectively. To guarantee balanced peak memory across devices, we add another two constraints, $\delta_{F}^{0}=\delta_{B}^{1}=\delta^{0}$ and $\delta_{F}^{1}=\delta_{B}^{0}=\delta^{1}$, where we use notations $\delta^{0}$ and $\delta^{1}$ for simplicity. As in Figure 3(d), we set $\delta^{0}=\delta_{F}^{0}=\delta_{B}^{1}=4$ and $\delta^{1}=\delta_{F}^{1}=\delta_{B}^{0}=2$. Note that we only control the offsets across different devices. For those adjacent passes within the same device, e.g. F and B of the last stage, B and W for all stages, we leave them flexible and only ensure their offsets are small (less than the repeating interval). According to Equation 1, we can analyse the asymptotic peak memory with respect to $d$,
peak memory of device $\text{peak memory of device }i\leq\frac{2d(\delta^{0}+\delta^{1})}{6}m+O(1)\approx\frac{\delta^{0}+\delta^{1}}{6}M$ (2)
By ignoring the small constant, we can directly control the peak memory by the value of $\delta^{0}$ and $\delta^{1}$.
3.2 V-Shape Pipeline Schedules
By varying the values of $\delta^{0}$ and $\delta^{1}$, we come up with 3 novel V-Shape building blocks (Figure 3), and present their final schedules based on our framework in Figure 4. The building block of V-Min (Figure 3(c)) has the minimum offsets, namely $\delta^{0}=\delta^{1}=1$, thus the minimum memory consumption. With $\delta^{0}=4$ and $\delta^{1}=2$ as in Figure 3(d), V-ZB eliminates the bubble to almost 0 (Figure 4(d)), pushing to extreme throughput. The building block of V-Half (Figure 3(b)), which uses $\delta^{0}=2$ and $\delta^{1}=1$, sits between the two extremes and consumes half of the activation memory required by 1F1B. Although both V-Min and V-Half have lower memory footprint than 1F1B, V-Min contains about 2/3 and V-Half contains about 1/2 of 1F1B’s bubbles, assuming F, B, W have equal run time. We show the comparison between our proposed V-Shape schedules and 1F1B in Table 1.
Table 1: For peak memory and bubbles, we ignore those small constant values. We redefine $\delta^{0}$/$\delta^{1}$ as the offsets between adjacent forward/backward passes for 1F1B.
Building Block $\delta^{0}$ $\delta^{1}$ Peak Memory Bubbles
1F1B $2$ $4$ $M$ $6d$
V-Min $1$ $1$ $M/3$ $4d$
V-Half $2$ $1$ $M/2$ $3d$
V-ZB $4$ $2$ $M$ 00
Notably, to avoid collision in building block, the offsets for V-Min and V-Half is slightly different on different $d$. The details are in Appendix F.
3.3 Repeating Bubbles in V-Min
In real cases where F, B and W have different run times, V-Min suffers from a repeating bubble. As shown in Figure 5, there exists a bubble for every repeating interval $T$. As a result, the bubble grows with respect to the number of microbatches. Although V-Half may also encounter the same issue, we surprisingly find that it generates patterns that tessellates well in most empirical cases. As illustrated in Figure 5(b), the throughput of V-Half is robust to the variation of run times. Additionally, we show that the bubbles of V-ZB will never grow when increasing the number of microbatches. We leave the related discussions in Appendix E.
@@ 3.3 Repeating Bubbles in V-Min
Figure 5: Repeating bubble in V-Min compared to V-Half. We take a repeating $d\times T$ grid from both schedule and assign F B W with different values. The result shows V-Min has a bubble for every repeating grid, while V-Half does not.
@@ 3.4 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks (see Appendix I), to show the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule with the same bubble rate as Interleaved 1F1B(Narayanan et al., 2021) but lower memory consumption.
@@ 4 Experiments
We construct our experiments to show three conclusions: a) The throughput and memory of V-Min, V-Half and V-Min aligns with the theoretical analysis in Section 3.2; b) Memory-saving methods including V-Min and V-ZB can bring accelerations; c) Our methods still perform best when combining with other state-of-the-art techniques.
@@ 4.1 Setup
We evaluate our methods using a serious of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 4 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stages to compensate for the embedding and output layer in LM, which can otherwise become the bottleneck of the pipeline and interfere the efficiency.
Table 2: Models used in experiments.
@@
Our experiments majorly focuses on the following pipeline parallel schedules: a) V-Min, V-Half and V-ZB: schedules introduced in Section 3.2; b) 1F1B and Interleaved 1F1B: methods implemented in Megatron-LM; c) 1F1B-R: 1F1B with full activation rematerialization (Chen et al., 2016); d) ZB-1P and ZB-2P: the adaptive zero-bubble methods introduced in (Qi et al., 2023) with activation memory limit set to the 1x/2x times of 1F1B.
@@ 4.2 Comparing Pipeline Schedules
In Figure 6, we present comparisons of the throughput measured in FLOPS utilization (MFU) and activation memory consumption across different pipeline schedules under various settings. From the results, V-ZB outperforms all other methods in terms of throughput, which aligns with Figure 4. When comparing the activation memory consumption, V-Min and V-Half stand out by significantly reducing activation memory to approximately 1/3 and 1/2, while other methods’ memory is similar. More details of our experiments and definition of metrics can be found in Appendix D.1.
Notably V-Min has a comparable throughput against 1F1B, but its throughput falls behind 1F1B at a larger number of microbatches due to the aforementioned repeating bubble in Figure 5(a), as discussed in Section 3.3. However, it still outperforms 1F1B with full activation rematerialization, providing a strong alternative for saving memory.
@@ 4.3 When to Save Memory
Another benefit of saving memory is that we can potentially use the extra memory for an increased microbatch size, which leads to a higher arithmetic intensity. We present the results in Figure 8. On bigger models, where memory pressure is higher and hence microbatch size is smaller, V-Half schedule can surpass V-ZB and other baselines because of its arithmetic intensity gain. This observation does not apply for V-Min, implying its arithmetic intensity gain can not compensate for the increased bubble. Doubling/Tripling the microbatch size for V-Half/V-Min results in a slightly higher activation memory than the other methods. This reflects the constant factors we ignored in Section 3.1. The increase is less significant as the number of devices grows.
@@ 4.4 Combining with Existing Techniques
We present our methods in the context of LLM training together with various other techniques. The following techniques are considered: a) Flash Attention (Dao et al., 2022; Dao, 2023); b) Tensor Parallelism (Narayanan et al., 2021) and Sequence Parallelism (Korthikanti et al., 2023); c) Distributed Optimizer provided in Megatron-LM. The implementations are all from Megatron-LM (Narayanan et al., 2021). Both our methods and the baseline methods are combined with the above techniques. Similar to the evaluation method in Kim et al. (2023), we perform a grid search on the following parameters: the size of PP; the size of TP; the size of DP; the microbatch size ($mbs$). We use 40 GPUs in this experiment. For each method, the best result from the grid search is reported.
We present the best result for each pipeline parallel schedule in Table 3 and the corresponding parameters. We find that when sequence length is smaller and hence the memory budget is more abundant, V-ZB performs the best due to the elimination of bubbles. When we increase the sequence length and the number of microbatches, V-Half performs the best because of its memory efficiency. The detailed data and analysis of grid search can be found in the Appendix D.3.
@@ 5 Conclusion And Future Work
In this work, we present a framework that constructs pipeline parallel schedules by focusing on their repeating building blocks. This framework enables direct computation of peak memory from the lifespan of the building block. Based on this capability, we design a family of memory-efficient building blocks. We discuss three representative methods from this family, namely V-Min, V-Half and V-ZB, and demonstrate with experiments that our methods advance the pareto frontier of throughput and memory in large model training. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules in future works.
In the future, we plan to further explore more memory efficient pipeline schedules based on our framework. A major limitation of V-Min is that, it suffers from growing bubbles when increasing the number of microbatches. Although V-Half mitigates this issue, there is still a space to further reduce the memory consumption. Using continuous offsets or finer-granularity discretization is a possible way to solve it. We leave it in our future work.
@@ Appendix I Other Memory Efficient Building Blocks
(b) Zero Bubble Schedule with 2/3 1F1B memory